OpenStudy (lgbasallote):

LGBADERIVATIVE: \[\huge \frac{d}{d\theta} \sec \theta \tan \theta\] i just use product rule right?

OpenStudy (lgbasallote):

\[\sec \theta (\sec^2 \theta) + \tan \theta (\sec \theta \tan \theta)\]

OpenStudy (lgbasallote):

\[\sec^3 \theta + \sec \theta \tan ^2 \theta\] does that sound right?

OpenStudy (kropot72):

It not only sounds right, it is right :)

OpenStudy (lgbasallote):

yay thanks!

OpenStudy (anonymous):

Yep! You get a medal :-3

OpenStudy (lgbasallote):

yayyy

OpenStudy (anonymous):

Want to try mine for a spin? ;-3 (active question)